Q2. (2 marks) 8 channels each with 25 kHz bandwidth are to be multiplexed together. In order to prevent interference there is a need for a guard band of 5 kHz between adjacent channels. How much bandwidth is needed to support the 8 channels?
Answer #1

In order to prevent interference, 7 guard bands are required for 8 channels.

Bandwidth required to support 8 channels = 8*25+ 7*5 = 200+35 = 235 kHz